OpenStudy (anonymous):

I have a trig test tomorrow, and I'm so stuck on a lot of these. Please help me out with this one, thanks! 6cosx +7tanx= secx

OpenStudy (ash2326):

\[6 cosx +7 tanx= secx \] \[6 \cos x+ 7 sinx/\cos x= 1/ \cos x\]

OpenStudy (ash2326):

\[6 \cos^2x+7 sinx =1\] we know cos^2 x+ sin^2 x=1 so cos^2 x = 1- sin^2 x substituting this in given equation 6(1-sin2^x)+7 sin x=1 6 - 6 sin^2 x+7 sin x=1 -6 sin^2 x+ 7 sin x+5=0 6 sin^2 x- 7 sin x-5=0 we know the standard quadratic equation ax^2+bx +c=0 x is given by quadratic formula \[x=(-b \pm \sqrt (b^2-4ac))/2a\] here \[sinx= (7 \pm \sqrt( 49-4*6*(-5))/12\] \[\sin x= (7 \pm \sqrt 169)/12\] \[\sin x= (7 \pm 13)/12\] sin x can't be 20/12 ( sin x<=1) sin x=-6/12 sin x= -1/2 \[x= 2npi -\pi/6\] \[x=( 2n+1) \pi + \pi/6\] n is an integer

OpenStudy (ash2326):

x= 7pi/6, 11pi/6, - pi/6 and so on

OpenStudy (ash2326):

did you get it

OpenStudy (anonymous):

I just don't understand how you got [6 \cos^2x+7 sinx =1\], the rest makes sense to me.

OpenStudy (ash2326):

we have 6cosx+7tanx=secx sec x can be written as 1/ cos x and tan x can be written as sinx/ cos x so 6cosx+7sinx/cosx=1/cosx now multiply both sides by cos x 6 cos^2x+ 7 sinx=1

OpenStudy (anonymous):

Ohhhhh ok, thank you so much :)

OpenStudy (ash2326):

welcome liz:)